The futures price of corn is $3.10 a bushel. Futures contracts for corn are based on 9,000 bushels, and the margin requirement is $2,000 a contract. You expect the price of corn to fall and sell the contract short.

  1. What is the value of the contract and how much must you initially remit? Round your answers to the nearest dollar.

    Value of the contract: $ _______

    Remit amount: $ _________

  2. If the futures price of corn rises to $3.19, what is the profit or loss on your position? Round your answer to the nearest dollar. Enter your answer as a positive value.

    The ___profit / loss___ is $ ________

  3. If the futures price of corn declines to $2.93, what is the profit or loss on your position? Round your answer to the nearest dollar. Enter your answer as a positive value.

    The ___profit / loss____ is $ _______

  4. If the futures price declines to $2.78, what may you do? Round your answer to the nearest dollar.

    The investor may remove up to $_____ from the account.

  5. How do you close your position?

    The investor closes the position by entering a contract to ___buy / sell___ the commodity.
Answer #1

Given Data

Future Price of Corn = $3.10
Contract Size = 9000 bushels
Margin Requirement = $2000
Position = Short (which means you are expecting that the price of corn will fall and you have sold the contract)

A.
We have to find
1. Value of Contract = No of bushels * Price of Corn
= 9000 * 3.10
Value of Contract = 27,900
2. Remit amount = Required Margin = $ 2000 = Given

B.
Price Goes Up to $ 3.19
As Price has gone up and we were expecting it to come down, which means we will incur a loss

Loss per Corn = Sell Price - Buy Price
  = 3.10 - 3.19
= 0.09
Loss (in$) = Loss per corn * No of bushels
  = 0.09 * 9000
Loss (in$) = 810

C.

Price Goes Down to $ 2.93
As Price has gone down and we were also expecting it to come down, which means we will incur a Profit

Profit per Corn = Sell Price - Buy Price
  = 3.10 - 2.93
= 0.17
Gain (in$) = Profit per corn * No of bushels
  = 0.17 * 9000
Gain (in$) = 1530

D.

Price Goes Down to $ 2.78
As Price has gone down and we were also expecting it to come down, which means we will incur a Profit

Profit per Corn = Sell Price - Buy Price
  = 3.10 - 2.78
= 0.32
Gain (in$) = Profit per corn * No of bushels
  = 0.32 * 9000
Gain (in$) = 2880
Investors can Remove Up To 2880

E.

The investor closes the position by entering a contract to buy the commodity

​​​​​​​Buy 9000 bushels contract